# Safe RLHF-V: Safe Reinforcement Learning from Human Feedback in Multimodal Large Language Models

### v1

@@ 5 Experiments
  • Q1: How does Safe RLHF-V compare to existing multimodal alignment algorithms, particularly in its ability to balance helpfulness and safety effectively?
  • Q2: Given that Safe RLHF-V relies on multimodal reward and cost models, how robust are these models with respect to preference data?
@@ 5.1 Experiment Setup
Datasets and Models To ensure the safety of MLLMs, we use Beavertails-V (in Section 3) as the training dataset. Specifically, we utilize the helpfulness and safety preference to train the RM-V and CM-V, respectively. Our experiments are conducted on three generic MLLMs: Llava-7B (version 1.5 and 1.6) and Qwen2-VL-7B. Furthermore, we fine-tune the original model using RLHF and DPO with single-dimension annotations and compare them against RLHF-V and Llava-RLHF as the comparison baseline.
Evaluation Metrics
Given the absence of publicly available multimodal evaluation datasets that simultaneously assess helpfulness and safety, we constructed our own evaluation set of BeaverTails-V. We employ the win rate metric to quantify improvements in model performance. Model outputs are evaluated using GPT-4o, and the specific prompts utilized for evaluation are detailed in Appendix.
@@
5.2 Main Results
To evaluate Safe RLHF-V’s ability to balance helpfulness and safety, we employ win rate as key metrics: We compute the win rate on the dedicated eval set of BeaverTails-V, where pairwise comparisons of model outputs (assessed via GPT-4o) indicate overall improvement.
As shown in Tab. 4, Safe RLHF-V delivers the best performance in balancing helpfulness and safety – optimizing helpfulness while ensuring compliance with safety constraints. Notably, Safe RLHF-V surpasses RLHF in safety and outperforms RLHF trained solely with helpfulness in the single dimension. For instance, in Beavertails-V, Safe RLHF-V consistently exceeds RLHF across various models in dual dimensions. A key factor behind this improvement is CM-V, which effectively constrains the exploration space during policy optimization, ensuring stable convergence.
Leveraging contrastive learning mechanism, single-dimensional DPO achieves remarkably high helpfulness scores. However, its performance in the safety dimension is considerably lower, highlighting the challenge of balancing helpfulness and harmlessness when RLHF does not explicitly model safety constraints. In real-world scenarios, preference data collected from diverse annotators inherently encapsulates multi-scale information due to variations in annotators’ backgrounds and values. Consequently, binary preferences alone often fail to ensure stable convergence, particularly in safety-critical settings.
5.3 Ablation Study
How much preference data do RM-V and CM-V require?
RM-V and CM-V serve as essential optimization signals within the Safe RLHF-V pipeline. To assess model accuracy across different data scales, we randomly selected varying proportions of preference data to train RM-V and CM-V separately. Our results indicate that the multimodal reward model achieves 82.2% accuracy with just 5K preference data points. In contrast, due to the complexity of safety constraints, CM-V attains only 54.5% accuracy under the same 5K safety preference data, highlighting the inherent challenges in modeling safety constraints. However, as the dataset size increases, CM-V’s accuracy improves, reaching 79% at 15K data points. From a practical perspective, this data volume remains acceptable for real-world applications.
Figure 4: The helpfulness and safety win rates of static reward shaping methods and Safe RLHF-V. Based on the Qwen2-VL-7B model, we find that static reward shaping coefficients tend to favor certain dimensions as their values increase, while Safe RLHF-V coefficients maintain a consistently high level. w/o and w denote the static reward shaping coefficients and Safe RLHF-V.
Effect of initial value of $\lambda_{0}$ with dynamic update and consistency update.
Under the min-max optimization, the Lagrange multiplier regulates the balance between RM-V and CM-V. We conducted comparative experiments by deploying a dynamic update mechanism and varying initial values of the multiplier on the Qwen2-VL-7B model. Using the dynamic update mechanism leads to the dynamic reward shaping coefficients, while remaining the fixed multiplier denotes the static ones (as shown in Fig. 4).
@@ 5.3 Ablation Study
From the results, when the multiplier remains fixed, the model’s performance deteriorates significantly as the initial value changes. The win rates of helpfulness and safety are formed like a parabola. This means that the model is very sensitive to the initial value of $\lambda_{0}$ and can hardly meet both good helpfulness and safety without carefully-assigned initial value. In extreme cases, the model results in collapse, or produce illogical outputs and even gibberish. On the contrary, we find when applying a dynamic update mechanism (Eq. 7), the final converged model’s helpfulness and safety show slightly fluctuations at a win rate of about 70%. This demonstrates the helpfulness and safety are both insensitive to the initial values at a high level.
5.4 The Training Curve and Analysis
As shown in Fig. 5, we visualize the evolution of $\lambda$ during training and its relationship with reward and cost, providing a clearer understanding of the constrained min-max optimization principle underlying Safe RLHF-V.
Figure 5: The trends of $\lambda$, reward, and cost reflect the constrained min-max optimization principle behind algorithm.
At the early stage of training, the model is influenced by factors such as toxicity and noise in the original training corpus, causing the cost to exceed the predefined threshold significantly. At this point, $\lambda$ is rapidly activated and quickly rises to its upper bound. During the subsequent period, $\lambda$ remains in a highly active state (upper bound), imposing strong penalties on the generation of unsafe responses. This gradually reduces the probability of the model producing unsafe outputs. In the oscillation phase, following the min-max optimization principle within a primal-dual optimization framework, $\lambda$ exhibits oscillatory behavior, accompanied by fluctuations in both reward and cost. During this phase, the cost hovers around the threshold as the model seeks an optimal solution that maximizes reward while satisfying the given constraints.
@@ 5.4 The Training Curve and Analysis
In the final convergence phase, both the cost and reward curves stabilize, while $\lambda$ gradually transitions into an inactive state (declining from the upper bound to zero). The red-shaded region in the figure illustrates that as the model undergoes fine-tuning, the probability of generating unsafe responses continues to decrease. At this point, the model is able to optimize its responses while adhering to safety constraints, significantly enhancing the safety of the reward model. This dynamic adjustment mechanism enables Safe RLHF-V to effectively improve model performance while maintaining compliance with safety constraints.
@@ 6 Conclusion and Outlook
In this work, we present Safe RLHF-V, a novel framework for aligning MLLMs with preference by optimizing for both helpfulness and safety. Our approach decouples the dual objectives into separate reward and cost models within a constrained MDP framework, thereby mitigating the trade-offs that traditional alignment methods fails. We also proposed the BeaverTails-V dataset, featuring dual preference annotations and multi-level harm severity labels. Furthermore, we introduce the multi-level safety guardrail model, Beaver-Guard-V, which bolsters the model’s resilience against toxicity queries. Experimental evaluations confirm that Safe RLHF-V yields significant improvements in both safety and helpfulness. Future work will extend this framework to incorporate additional modalities and adaptation mechanisms to address evolving safety challenges in AI systems.
Author Contribution
J.J initiated the project, the manuscript writing, and experimental design. X.C. developed the initial version of the Safe RLHF codebase and conducted the experiments on LLaVA 1.5 7B. R.P, J.J and X.C curated and analyzed the datasets. R.P further evaluated the performance of Safe RLHF on Qwen 2.5 and conducted experiments on the Guard Model. H.Z carried out the experiments on LLaVA 1.6. C.Z contributed to experimental discussions and data visualization. J.L assisted in dataset categorization and taxonomy design. D.H, B.C, J.Z, K.W, J.D, and C.C participated in experimental design, discussions, and code verification. S.H and Y.G provided valuable suggestions on writing and ablation studies. Y.Y led the project and supervised the overall progress.
@@ A.1 Preference Models
We initialize our reward model (RM-V) and cost model (CM-V) using the pre-trained model, Llava-1.5-7B [15]. During the training phase, we employ the loss functions presented in equations (4) and (4.2). Additionally, we incorporate an extra regularization term within the loss function to enhance generalization and stabilize the training process.
@@ A.2 Details of RLHF Training
Following the training paradigm proposed by Ouyang et al. (2022), we use reinforcement learning from human feedback (RLHF) to optimize our model. The training objective consists of two key components: the RL objective and the PTX pretraining objective. The RL objective is guided by a reward model-vision (RM-V), with an additional per-token KL penalty to constrain policy updates and ensure stable learning.
@@
For RLHF fine-tuning, we adopt the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017), employing a clipped surrogate loss formulation:
@@
where $\theta_{\text{old}}$ represents the model parameters from the previous update, and $\lambda\in(0,1)$ is the PPO clipping coefficient. The advantage estimate $A_{t}$ is computed using Generalized Advantage Estimation (GAE) (Schulman et al., 2015).
@@
We utilize the Align-Anything-TI2T-Instruction-100K Dataset (Ji et al., 2024c) for PTX optimization. The total training loss during the RLHF phase is defined as follows:
@@ A.3 Details of Safe RLHF-V Training
Similar to the Safe RLHF training process proposed by Dai et al. (2024), Safe RLHF-V iteratively solves the minimax problem in equation (7) by alternately updating the model parameters $\theta$ and the Lagrange multipliers $\lambda$.
@@
Appendix B Detailed Description of Meta Labels

### v2

@@ 5 Experiments
  • Q1: How does Safe RLHF-V compare to existing multimodal alignment methods?
  • Q2: How robust are reward and cost models in Safe RLHF-V with respect to preference data?
@@ 5.1 Experiment Setup
Datasets and Models.  To ensure the safety of MLLMs, we use Beavertails-V as the training dataset. Specifically, we utilize the helpfulness and safety preference to train the RM-V and CM-V, respectively. Our experiments are conducted on various MLLMs: Llava-7B-(1.5 and 1.6) and Qwen2-VL-7B. Furthermore, we fine-tune the original model using RLHF [6] and DPO [7] with single-dimension annotations and compare them against RLHF-V [11], Llava-RLHF [12], and MM-RLHF [49] as the comparison baseline.
Evaluation Metrics.  Given the absence of publicly available multimodal evaluation datasets that simultaneously assess helpfulness and safety, we constructed our own evaluation set of BeaverTails-V. We employ the win rate metric to quantify improvements in model performance. Model outputs are evaluated using GPT-4o, and the specific prompts utilized for evaluation are detailed in Appendix C.
5.2 Main Results
We use win rate as the key metric, and the pairwise comparisons of model outputs (assessed via GPT-4o) indicate overall improvement. As shown in Table 3, Safe RLHF-V delivers the best performance in optimizing helpfulness while ensuring compliance with safety constraints. Notably, Safe RLHF-V surpasses RLHF in safety and outperforms RLHF trained solely with helpfulness in the single dimension. For instance, in Beavertails-V, Safe RLHF-V consistently exceeds RLHF across various models in dual dimensions. A key factor behind this improvement is CM-V, which effectively constrains the exploration space during policy optimization, ensuring stable convergence. While single-dimensional DPO achieves high helpfulness scores via contrastive learning, its safety performance is significantly lower, underscoring the difficulty of balancing helpfulness and harmlessness without explicit safety modeling. In real-world scenarios, preference data from diverse annotators reflects varying values and backgrounds, introducing multi-scale information. As a result, binary preferences alone often fail to ensure stable convergence, especially in safety-critical settings.
Figure 3: Left: Ablation: static reward shaping and Safe RLHF-V. We find that reward shaping tends to improve single dimension, while our method maintains a consistent improvement for dual objective. w/o and w denote the static reward shaping coefficients and Safe RLHF-V. Right: The trends of $\lambda$, reward, and cost reflect the constrained min-max optimization principle behind the algorithm.
5.3 Ablation Study
How much preference data do RM-V and CM-V require?
RM-V and CM-V serve as essential optimization signals within the Safe RLHF-V pipeline. To assess model accuracy across different data scales, we randomly selected varying proportions of preference data to train RM-V and CM-V separately. As shown in Table 4, our results indicate that the multimodal reward model achieves 82.2% accuracy with just 5K preference data points. In contrast, due to the complexity of safety constraints, CM-V attains only 54.5% accuracy under the same 5K safety preference data, highlighting the inherent challenges in modeling safety constraints. However, as the dataset size increases, CM-V’s accuracy improves, reaching 79% at 15K data points. From a practical perspective, this data volume remains acceptable for real-world applications.
@@ 5.3 Ablation Study
Sensitive of initial value of $\lambda_{0}$ We conducted the comparison between dynamic update mechanism and reward shaping with different initial values of the multiplier on the Qwen2-VL-7B model. As shown in Fig. 3, when the multiplier remains fixed, the model’s performance deteriorates significantly as the initial value changes. The win rates of helpfulness and safety are formed like a parabola. This means that the model is very sensitive to the initial value of $\lambda_{0}$ and can hardly meet both good helpfulness and safety without a carefully-assigned initial value. In extreme cases, the model results in collapse or produces illogical outputs and even gibberish. On the contrary, we find that when applying a dynamic update mechanism (Eq. 7), the final converged model’s helpfulness and safety show slight fluctuations at a win rate of about 70%. This demonstrates that the dynamic update mechanism is insensitive to the initial values.
5.4 The Training Curve and Budget Bound Analysis
As shown in Fig. 3, we illustrate the evolution of $\lambda$ during training and its correlation with reward and cost, offering clearer insights into the constrained min-max optimization principle of Safe RLHF-V. In the early training stage, the unaligned initial model exceeds the predefined safety threshold. At this stage, $\lambda$ is rapidly activated and quickly reaches the budget bound. Subsequently, $\lambda$ stays at its upper bound, strongly penalizing the generation of unsafe responses. This gradually lowers the model’s likelihood of generating unsafe outputs. In the oscillation phase, $\lambda$ follows the min-max principle, exhibiting oscillations along with reward and cost fluctuations. During this phase, the cost remains near the threshold as the model seeks to maximize reward under the given constraints.
@@ 5.4 The Training Curve and Budget Bound Analysis
In the final convergence phase, the cost and reward curves stabilize, and $\lambda$ gradually becomes inactive, declining from budget bound to zero. The red-shaded area shows that during fine-tuning, the model’s probability of generating unsafe responses steadily decreases. At this stage, the model optimizes responses while adhering to safety constraints, greatly improving reward model safety. This dynamic adjustment enables Safe RLHF-V to enhance model performance while ensuring safety compliance.
@@ 6 Conclusion and Outlook
This work introduces Safe RLHF-V, the first comprehensive framework for multimodal safety alignment. Central to our approach is BeaverTails-V, the first open-source dataset with dual preference annotations and multi-level safety labels, enabling fine-grained supervision of both helpfulness and safety. Through constrained optimization and the Beaver-Guard-V defense system, our method achieves significant improvements in safety (+34.2%) and helpfulness (+34.3%). These results highlight the potential of dual-supervised data and principled optimization for building safer, more capable AI assistants. While our study focuses on image-text scenarios, broader generalization to other modalities remains a limitation. Future work will extend this framework to incorporate additional modalities and adaptation mechanisms to address evolving safety challenges in AI systems.
@@ A.1 Preference Models
We initialize our reward model (RM-V) and cost model (CM-V) using the pre-trained model, Llava-1.5-7B [3]. During the training phase, we employ the loss functions presented in equations (4) and (5). Additionally, we incorporate an extra regularization term within the loss function to enhance generalization and stabilize the training process.
@@ A.2 Details of RLHF Training
Following the training paradigm proposed by [6] ([6]), we use reinforcement learning from human feedback (RLHF) to optimize our model. The training objective consists of two key components: the RL objective and the PTX pretraining objective. The RL objective is guided by a reward model-vision (RM-V), with an additional per-token KL penalty to constrain policy updates and ensure stable learning.
@@
For RLHF fine-tuning, we adopt the Proximal Policy Optimization (PPO) algorithm ([31]), employing a clipped surrogate loss formulation:
@@
where $\theta_{\text{old}}$ represents the model parameters from the previous update, and $\lambda\in(0,1)$ is the PPO clipping coefficient. The advantage estimate $A_{t}$ is computed using Generalized Advantage Estimation (GAE) ([50]).
@@
We utilize the Align-Anything-TI2T-Instruction-100K Dataset ([51]) for PTX optimization. The total training loss during the RLHF phase is defined as follows:
@@ A.3 Details of Safe RLHF-V Training
Similar to the Safe RLHF training process proposed by [16], Safe RLHF-V iteratively solves the minimax problem in equation (7) by alternately updating the model parameters $\theta$ and the Lagrange multipliers $\lambda$.
@@
Appendix B Annotation Pipeline
@@ B.1 The Pipeline of Data Generation and Annotation
Image Collection:
Leveraging our defined harm taxonomy, we used the Yandex search engine as the primary source for diverse image retrieval. GPT-4 generated category-aligned keywords, which were manually curated and refined to minimize discrepancies between retrieved images and category expectations. To balance representation across harm levels, we dynamically adjusted search strategies, intensifying retrieval for high-harm categories. All collected images underwent multiple rounds of manual filtering and validation to ensure quality and relevance.
Prompt Generation:
Generating prompts that are both meaningful and contextually relevant to the images is crucial for capturing the complexity of harmful content. To address this, we combined instruction-following and jailbreak strategies. Using the Ovis1.6-Gemma2-9B model, we generated eight prompt variations per image, striking a balance between diversity and quality.
@@ B.2 Harm Classification (may contain harmful examples)
Figure 5: Example of harmful and unethical themes in BeaverTails-V: A collection of 20 panels categorized into 9 primary harmful domains, each labeled with a number (X-X), where the first X indicates the primary category and the second X represents the subcategory, covering a range of harmful and unethical behaviors.